Using the financial statements for the Snider Corporation, calculate the 13 basic ratios found in the chapter. SNIDER CORPORATION Balance Sheet December 31, 20X1 Assets Current assets: Cash $ 57,100 Marketable securities 26,100 Accounts receivable (net) 184,000 Inventory 242,000 Total current assets $ 509,200 Investments 64,500 Plant and equipment. $689,000 Less: Accumulated depreciation 256,000 Net plant and equipment 433,000 Total assets $ 1,006,700 Liabilities and Stockholders' Equity Current liabilities: Accounts payable $ 92,700 Notes payable 76,500 Accrued taxes 17,000 Total current liabilities $ 186,200 Long-term liabilities: Bonds payable 151,000 Total liabilities $ 337,200 Stockholders' equity Preferred stock, $50 par value $ 100,000 Common stock, $1 par value 80,000 Capital paid in excess of par 190,000 Retained earnings 299,500 Total stockholders' equity $ 669,500 Total liabilities and stockholders' equity $ 1,006,700 SNIDER CORPORATION Income Statement For the Year Ending December 31, 20X1 Sales (on credit) $ 2,076,000 Cost of goods sold 1,326,000 Gross profit $ 750,000 Selling and administrative expenses 490,000 * Operating profit (EBIT) $ 260,000 Interest expense 31,800 Earnings before taxes (EBT) $ 228,200 Taxes 84,400 Earnings after taxes (EAT) $ 143,800 *Includes $36,400 in lease payments. Using the above financial statements for the Snider Corporation, calculate the following ratios.

Profitability ratios. (Do not round intermediate calculations. Input your answers as a percent rounded to 2 decimal places for every question).
  

A. Profit Margin

Return On assets( Investment )

Return on equity

B. Receivable turnover

Average collection period(days)

Inventory Turn Over

Fixed asset turnover

Total asset turnover

C. Current ratio

Quick ratio

D. Debt to total assets

Times interest earned

Fixed coverage charge
Answer #1

As per rules I am answering the first 4 subparts of the question

1. Profit Margin= Net Income/ Revenue

= 143800/2076000

=6.93%

2: Return On assets( Investment )= Net Income/ Total assets

= 143800/1006700

=14.28%

3 Return on equity = Net Income available to equity shareholders/ Equity

= 143800/ 669500

=21.48%

4: Receivable turnover = Credit sales/ Accounts receivable

= 2076000/184000

=11.28